### v1

# Exoplanets synchronization: Learning from Venus’ retrograde rotation

Abstract
Context: Planets in the HZ can have dense atmospheres affecting their rotations. Over time, the rotation tends to stationary solutions that can be synchronous or asynchronous.
Aims: Our understanding of Venus’s rotational dynamics is revisited to look at what might happen to exoplanets in the habitable zone of a solar-type star.
Methods. The creep tide theory is used to calculate the gravitational tidal torque. Mathematical analysis is used to study the differential equation resulting from the joint contributions of tidal and atmospheric torques.
Results. The formation of a dense atmosphere can alter the primordial rotation of the planet. One possibility is that it gradually becomes retrograde. The rotation of Venus is an example.
keywords:
Exoplanets - Venus - Rotation
1 Introduction
This paper aims to use our knowledge of Venus’ rotation as an insight into what to expect for the rotation of Earth-like exoplanets in the habitable zone of a solar-type star. It collects various published results and compares them to hypothetical situations to improve our understanding of their meanings. It includes a new discussion of the existence and stability of stationary solutions under the joint actions of tidal and atmospheric torques (Auclair-Desrotour et al. 2017). When the eccentricity is neglected, these solutions might be synchronous or asynchronous (a.k.a. parasynchronous), stable or unstable (Correia and Laskar, 2003; Leconte et al., 2015). Two parameters indicating the ratios of the tidal and atmospheric torques are introduced, as well as a toy model. They allow us to study several different evolutionary paths for the rotation of exoplanets located in the habitable zone of solar-type stars. In these cases, the HZ is not too close to the star, and planets may harbor dynamically important atmospheres. The exploration of the model shows that the evolution leading to a retrograde rotation can happen at a slow pace and is not a low-probability event. The retrograde rotation of Venus may be the result of one smooth process of this kind, associated with the formation of its atmosphere, rather than a collisional one that has transferred a huge amount of negative angular momentum to the planet. The collisional processes often used to explain the retrograde rotation of Venus (McCord, 1968; Makarov and Goldin, 2023) are not the only possibility.
In section 2, the new full 3D version of the creep tide theory (Folonier et al. 2025) is used to study the evolution of the rotation of a hypothetical naked Venus in both the prograde and retrograde rotation cases. In particular, it is used to investigate how long the retrograde rotation can resist the action of the tidal torque. The precession of the planet’s equinoxes and the variations in the obliquity are also studied. In section 3, the torque due to the phase lead of the atmospheric bulge is introduced, and the main effects due to the joint actions of the atmospheric and tidal torques are investigated. In section 4, it is shown how the evolution of the atmosphere can cause the reversal of the planet’s rotation from prograde to retrograde. A toy model for the planet rotation provides evolutionary examples of this process. Section 5 collects the main conclusions of the report.
2 The naked Venus
Venus currently has a retrograde rotation with a period $243.0226\pm 0.0013$ days (Margot et al. 2021). We know that this rotation is strongly affected by the thermal bulges in the planet’s atmosphere. How would a retrograde rotation of this kind evolve should the planet be devoid of an atmosphere? We have investigated this topic using the new version of the creep tide theory valid for systems with a nonzero obliquity (Folonier et al., 2025).
2.1 Back to prograde
The main result is that the retrograde rotation of a naked Venus would be changed to prograde in a relatively short time (see Fig. 1). Using typical values for Venus’ parameters involved in this calculation (see Table 1), the result is an inversion in $\sim 700,000$ years, if the relaxation is $\gamma\sim 2\times 10^{-7}{\rm s}^{-1}$ (as the solid Earth cf. Ferraz-Mello, 2013). The tidal theory also indicates that at the inversion, we do not have a reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the completion of the inversion. During the inversion interval, the angular momentum of the planet is very small (the rotation period grows to several thousand days). We do not know if the planet itself would turn upside down because the equations of the polhode are not included in the model. After the reversal of the rotation back to a prograde state, the rotation accelerates and is driven to a stable synchronization with the orbital period (see Fig. 1 Bottom).
Figure 1: Variation of the rotational elements of the planet before and after the rotation reversal: Top and Middle: Obliquity and Equinox before and after the rotation reversal. Bottom: Rotation period and Orbital period (blue line). Longitude of the First Equinoctial Point (intersection of the planet’s equator and orbit)
Table 1: Venus Dynamic Parameters
Mass ($M$) 0.815 Earth mass
Radius ($R$) 0.949 Earth Radius
Rotation Period 243.0226 d (Retrograde)
Oblateness $6.64\times 10^{-6}$
Axial tilt with respect to the ecliptic pole ($J$) $175^{\circ}.7$
[…]
Moment of Inertia ($C$) 0.34 $MR^{2}$
Tidal Relaxation Factor ($\gamma$) $2\times 10^{-7}\ {\rm s}^{-1}$
Tidal Love number ($k_{2}$) 0.295
Semi-major axis 0.7238 AU
Orbital period 224.9145 d
Eccentricity 0.006773
Inclination $3^{\circ}.395$
Length of the day 116.8091 d
Atmospheric Bulge Lead ($\sigma^{\prime}$) $30-34^{\circ}$
Semi-diurnal frequency ($\nu)$ $1.245144\times 10^{-6}{\rm s}^{-1}$
[…]
2.2 Precession
The other striking result given by the creep tide theory is that the motion of the poles in the celestial sphere, the equinox precession, is not a circulation, as shown with rigid-body models. It is rather a libration whose amplitude decreases up to a short time before the inversion (See Fig. 1 Top). During inversion, the equinoctial points circulate, and afterward they again librate. In Fig. 2, one may note the reversal of the direction of circulation of the precession when obliquity crosses $\beta_{\rm pl}=90^{\circ}$ (dashed line).
In the calculations leading to the results shown in Fig. 1, we have fixed the oblateness of the planet at $6.64\times 10^{-6}$. It is known that the oblatenesses of slow-rotating bodies (as Mercury and Venus) are much larger than predicted by the usual Maclaurin or Clairaut formulas. The model adopted here opts for the introduction of the oblateness by hand when the rotation velocity falls below a certain lead-off limit. This is done following the recipe: If $T_{\rm rot}\geq 33.528\$days, then $\epsilon_{\rm M}=6.42\times 10^{-6}$. The chosen lead-off period is the rotation period corresponding to the oblateness calculated from the $J_{2}$ value determined in recent space missions. The difference $0.22\times 10^{-6}$ between the two values mentioned above comes from the fact that the actual oblateness of the body is the sum of the rotational component and one small component due to the tidal deformation: $\epsilon_{\rm Z}=\epsilon_{\rm M}+\frac{1}{2}\epsilon_{\rm Jeans}$ (see Ferraz-Mello, 2015).
Tests done using this model and the original creep tide model using the Maclaurin formula for the oblateness showed that the evolution of the non-angular parameters of the rotation and the orbital elements are not affected by this change. Only the angles fixing the planet’s equator are changed. The influence of oblateness on the precession is important. The application of the same model using the observed oblateness to calculate the solar component of the Earth’s precession gives for the period a result (79,476 yr) very close to the result obtained using classical rigid body models.
Figure 2: Motion of the pole in the interval $660<t<720$ (kyr). Polar coordinates: Obliquity and Longitude of the First Equinoctial Point ($\Omega_{\rm Ipl}$). Dashed line: $\beta=90^{\circ}$.
2.3 Prograde rotation dynamics
Let us assume that the primordial rotation of Venus was prograde. How would this rotation be modified by tides in the absence of an atmosphere? The simulations show very different behaviors for fast and slow rotating planets. Fig. 3 shows (magenta lines) the results obtained assuming a solid Earth-like tidal relaxation ($\gamma=2\times 10^{-7}{\rm s}^{-1}$), and an initial rotation period of 10 days (terrestrial days). The almost synchronous stationary solution is reached in about 107 Myr  [^1] The stationary solution is slightly supersynchronous by an amount of the order of $e^{2}$. In the case of terrestrial planets, $T_{\rm pl}\sim T_{\rm orb}(1-6\gamma^{2}e^{2}/n^{2})$ (see Ferraz-Mello, 2013; Ferraz-Mello et al., 2025) which gives for a prograde Venus, a difference of the two periods $\Delta T/T\sim 10^{-4}$. With a nonzero eccentricity, synchronization can only be reached if the almost synchronous rotation entails the outcrop of an asymmetry in the mass distribution of the planet the torque of which counterbalances the small torque surplus due to eccentricity. The obliquity of the planet grows monotonically, but before reaching the stationary solution, it has a maximum and is then driven to zero in a few million years. Note that no obvious symmetry appears in the simulation as a cause of this behavior. This behavior seems very robust and does not depend critically on the initial conditions. Simulations with small variations in the initial conditions show only small, regular quantitative differences (see fig. 4). The chaotic behavior predicted by Laskar and Robutel (1993) does not affect the solutions because the secular perturbations due to the other planets of the system are not being considered. The equinoctial nodes precess initially with a period of $\sim 130$ kyr and grow following a quadratic time law to reach about 320 kyr in 100 Myr. When obliquity is driven to zero, the equinoctial nodes are no longer defined, and the circulation is converted to libration after a period of somewhat irregular behavior.
Figure 3: Left: Tidal evolution of the obliquity and rotation period of a prograde naked Venus whose rotation period at t=0 is 10 days. Magenta: Case of a planet with the same tidal relaxation as the solid Earth ($\gamma=2\times 10^{-7}{\rm s}^{-1}$). Blue: A softer case ($\gamma=2\times 10^{-6}{\rm s}^{-1}$). Right: Zoom of the evolution in the neighborhood of the synchronization.
Figure 4: Same as fig. 3 for a cluster of close initial conditions.
2.4 The rotation phase diagram
In the zero-obliquity version (a.k.a. coplanar version) of the tidal theory, the variation of the rotation of the planet is given by a first-order differential equation (Ferraz-Mello, 2013; Correia et al., 2014) of the form
$\dot{\Omega}=-A\sum_{k\in Z}E^{2}_{2,k}\sin 2\sigma_{k}$ (1)
where $A$ is a positive coefficient depending on the dynamical parameters of the system, $E_{2,k}$ are known functions of the orbital eccentricity (Cayley polynomials), and $\sigma_{k}$ are angles related to the lags of the components of the tide, defined by
$\tan\sigma_{k}=\frac{\nu+kn}{\gamma}$ (2)
where $\nu=2\Omega-2n$ is the semi-diurnal frequency (Ferraz-Mello, 2013; Ferraz-Mello et al., 2025). Hence,
$\sin 2\sigma_{k}=\frac{2\gamma(\nu+kn)}{\gamma^{2}+(\nu+kn)^{2}}.$ (3)
In the circular approximation, the only remaining term is $k=0$ and $\sigma_{0}$ is the lag of the main tidal component: the semi-diurnal tide.
Figure 5: Top: Phase diagram for $e=0.006773$ and relaxation factors $2\times 10^{-7}{\rm s}^{-1}$ (solid line), $7\times 10^{-7}{\rm s}^{-1}$ and $2\times 10^{-6}{\rm s}^{-1}$ (dashed lines). Bottom: Same as top for $e=0.3$.
Figure 5 Top is the plot of the function $\dot{\Omega}=f(\Omega)$ for a wide range of rotation velocities of a naked Venus. It shows that for the current orbital eccentricity, there is only one stationary solution $\mathsf{S}$, at $\Omega/n\simeq 1$. This solution is stable since the derivative $\dot{\Omega}$ is positive on the left of $\mathsf{S}$ and negative on its right. This situation is very robust and is verified for a wide range of variations of the parameters involved in the function $f(\Omega)$. However, for large eccentricities, it changes. Figure 5 Bottom shows the results corresponding to e=0.3. In this case, the stationary solution is displaced to the point $\Omega/n=3/2$. The stationary solution now corresponds to a spin-orbit resonance of the same kind as that observed for Mercury. The lower figure is atypical. Generally, as seen for Mercury (Correia et al., 2014; Ferraz-Mello, 2015) and for several exoplanets, a stable stationary solution may exist at $\Omega/n=3/2$, but the almost synchronous stationary solution continues to be seen in the phase diagram. Here, the synchronous solution disappears when $e=0.3$. The possibility of an artifact resulting from sampling seems discarded. The plot remains as shown even with sampling steps as small as $\Delta(\Omega/n)=10^{-4}$.
3 The atmospheric torque
We have seen that the retrograde rotation of Venus could not survive for a long time, and the observed retrograde rotation is a clear clue to the existence of a torque acting in that direction. It has been attributed to a phase lead in the thermal deformation of the atmosphere (Gold and Soter, 1971; Ingersoll and Dobrovolskis, 1978)  [^2] For detailed discussions of this model, see Correia and Laskar(2003), Leconte et al.(2015), Auclair-Desrotour et al. (2017, 2024), and references therein.
Solar radiation raises the air temperatures in the afternoon and causes the air mass to flow to colder regions. In a first approximation, the atmosphere may be represented by an ellipsoid whose bulge is displaced in the direction contrary to the rotation of the body (see fig. 6). In this figure, the direction of the orbital motion does not matter. The rotation direction indicated is not absolute, but relative to the Sun’s direction. At one point on the planet’s surface, the tidal bulge is felt after noon. The tide lags the forcing tidal forces. In contrast, the atmospheric bulge leads the rotation. The gravitational effects of the Sun on the atmosphere introduce a torque in a direction opposite to that of the gravitational tidal torque.
Figure 6: The atmospheric and tidal bulges in a system of reference in which the direction of the Sun is kept fixed. cf. Ingersoll and Dobrovolskis (1978).
Using notations similar to those adopted in the previous Section, the contribution of the atmospheric bulge to the rotation of the planet is
$\dot{\Omega}=A^{\prime}\sin 2\sigma^{\prime}$ (4)
where $A^{\prime}$ is a positive coefficient  [^3] The coefficient $A^{\prime}$ depends on the mean motion ($n$), the radius ($R$), the moment of inertia ($C$), and also on the temperature at the surface, ($T_{0}$), on its specific heat at constant pressure, ($c_{p}$), and on the heat flux at the sub-solar point, ($\Phi_{0}$). It is such that $A^{\prime}\gamma^{\prime}=\frac{3\pi n^{2}R^{4}\Phi_{0}}{8c_{p}T_{0}C}.$ (5) This coefficient formally differs from the coefficient appearing in Ingersoll and Dobrovolskis(1978) by a factor 2. The division by 2 is here necessary to compensate the factofor 2 present in the expression of $\sin 2\sigma^{\prime}$. $A^{\prime}$ has the same dimension as $A$. and $\sigma^{\prime}$ is the lead of the atmospheric bulge defined by
$\tan\sigma^{\prime}=\frac{\nu}{\gamma^{\prime}}\equiv\nu\tau^{\prime}$ (6)
(Ingersoll and Dobrovolskis, 1978). Note that when $e=0$ equations (1) and (4) are formally equal. Since the given definitions of $\sigma$ and $\sigma^{\prime}$ are not univocal, we stress that in both cases they are defined by the principal value of the function, in the interval $(-\pi/2,+\pi/2)$. In the case of Venus, the maxima of the atmospheric pressure are recorded at 09:45 and 22:00 local solar time (see Auclair-Desrotour et al. 2024). This means $\sigma^{\prime}\sim 32^{\circ}$ .
3.1 Composite phase diagram
Under hypotheses of zero obliquity and zero eccentricity, the variation of the rotation of the planet is given by an autonomous first-order differential equation formed by the composition of Eqns. (1) (for $e=0$) and (4):
$\dot{\Omega}=-A\sin 2\sigma_{0}+A^{\prime}\sin 2\sigma^{\prime}$ (7)
or
$\dot{\Omega}=A\sin 2\sigma_{0}\left(b\ \frac{\nu^{2}+\gamma^{2}}{\nu^{2}+\gamma^{\prime 2}}-1\right)$ (8)
where
$b=\frac{A^{\prime}\gamma^{\prime}}{A\gamma}.$ (9)
The parameter $b$ is akin to the ratio of the two torques. $b=0$ in the absence of atmospheric torque and it increases with the influence of the atmospheric torque. It plays a critical role on the existence and stability of stationary solutions.
The red curve in Fig. 7 is the plot of the resulting function $\dot{\Omega}(\Omega)$ in arbitrary units (such that the largest elongation of the function is equal to 1). The figure also shows the individual contribution of the two terms of Eqn. (7): the tidal torque (black solid-line) and the atmospheric torque (dashed line).
The intersections of the red curve with the axis $\dot{\Omega}=0$ are stationary solutions of the system. One of these solutions, $\mathsf{S}$, lies at $\Omega/n=1$ and is a synchronous solution. A simple analysis of the sign of $\dot{\Omega}$ on the right and left of this point indicates that this solution is unstable. The derivative of the curve $\dot{\Omega}(\Omega)$ in the neighborhood of the crossing point is positive. One may note that in the component corresponding to the tide of a naked planet (black solid line), the sign is negative, indicating the stability of the synchronous solution in that case.
The figure shows two more intersections at the points $\mathsf{P}$ and $\mathsf{P}^{\prime}$. The derivative of the functions at these two crossings of the axis $\dot{\Omega}=0$ is negative, so both correspond to stable stationary solutions. These two asynchronous solutions are symmetric with respect to the synchronous solution and, in the case shown in Fig. 7, the solution at $\mathsf{P}$ is prograde and the solution at $\mathsf{P}^{\prime}$ is retrograde.
In the case of Venus, the figure is similar to Fig. 7. However, the excursion of the function $\dot{\Omega}(\Omega)$ near the synchronous solution is too large and, keeping the figure at scale, the red lines near the asymmetric solutions and the axis $\dot{\Omega}=0$ are indistinguishable, leaving the crossings invisible (even if they are well defined in the data file). To make them visible, the factors $\gamma$ and $\gamma^{\prime}$ adopted in the construction of Fig. 7 were multiplied by 10.
Figure 7: Phase diagram $\dot{\Omega}(\Omega)$ resulting from the composition (red) of the tidal (solid black) and atmospheric (dashed) torques. The intersections of the red curve with the axis $\dot{\Omega}=0$ (brown) are the stationary solutions: The unstable synchronous solution $\mathsf{S}$ and two stable asynchronous solutions $\mathsf{P}$ and $\mathsf{P}^{\prime}$, one prograde and one retrograde. The adopted factors $\gamma$ and $\gamma^{\prime}$ were exaggerated to improve the visibility of the plot’s main features.
3.2 Existence and stability of the asynchronous solutions
Equation (8) can be expressed as the quotient of two polynomials in $\nu$, and equation $\dot{\Omega}=0$ has the same roots as the symmetric third-degree polynomial in the numerator. One of these roots is $\nu_{0}=0$, and the other two are given by
$\nu_{0}^{2}=\gamma^{\prime 2}\ \frac{1-b\Gamma^{2}}{b-1}$ (10)
where
$\Gamma=\frac{\gamma}{\gamma^{\prime}}.$ (11)
The condition for the existence of real roots is that the numerator and denominator are not simultaneously positive or negative (in which case we get $\nu_{0}^{2}<0$ and the two roots are imaginary). The conditions for the existence of the real roots $\pm\nu_{0}$ are:
Case I
[…]
Case II
$\qquad 1<b<\Gamma^{-2}\qquad(\Gamma<1)$.
Otherwise, the two roots are imaginary.
The analysis of the behavior of $\dot{\Omega}(\Omega)$ when $|\Omega|\rightarrow\infty$ allows us to conclude that, in case I, the synchronous root is unstable and the two asynchronous roots are stable (as shown in Fig. 7). In case II, the behavior is inverted, the synchronous root becomes stable and the two asynchronous roots become unstable.
Figure 8: Loci of the asynchronous stationary solutions for given values of $\Gamma$. These solutions are stable if they fall in the zone $b<1$ and unstable otherwise. If $|\nu_{0}/n|<1$ they are both prograde. If $|\nu_{0}/n|>1$, one is prograde and the other is retrograde.
Figure 9: Domains of existence of the asynchronous stationary solutions and vice versa, i stable, II unstable. Outside these domains, they are not real. In domains 0 and II, the synchronous stationary solution is stable. In other domains, it is unstable. The inserted sketches show how the roots are formed in each domain.
These stationary solutions are similar to those found by Auclair-Desrotour et al. (2017) using the classical Darwin tidal theory with Maxwell rheology.
3.3 Stability of the synchronous solutions
In domains I and II the synchronous solution is unstable when the asynchronous solutions are stable, and vice versa. In the other domains, the synchronous solution is the only real root. The stability is again deduced from the study of the leading term of the third-degree polynomial that defines the roots: $\dot{\Omega}\propto 2B(b-1)\nu_{0}^{3}+\&c.$ [^4] The function is of the kind y’=y.f(y). The stability of the root $y=0$ may be inferred from the study of a finite interval. The synchronous solution is stable when $b<1$ and unstable otherwise. This is consistent with the plots of the partial contributions of the tidal and atmospheric torques shown in Fig. 7.
In the domain 0 (green area) of Fig. 9, the tidal torque is dominant and the synchronous solution is stable as in the case of a naked Venus. In the white area, the atmospheric torque is dominant, and the synchronous solution is unstable.
This result depends on the theory of tidal friction used. The Darwin theory with a Maxwell rheology (Auclair-Desrotour et al. 2017) gives results equal to those discussed above. However, if a constant lag or the lag given by Andrade’s rheology is plugged into Darwin’s theory, the asynchronous solutions come in pairs (Leconte et al. 2015), and the synchronous solution remains stable, even in the presence of stable asynchronous solutions.
4 The evolution towards a retrograde rotation
In the general case, planets initially have a fast prograde rotation. Figure 7 shows that in case I, it is impossible for the rotation to evolve smoothly to a retrograde condition. The space between the parasynchronous solutions $\mathsf{P}$ and $\mathsf{P}^{\prime}$ acts as a wall that prevents one solution starting on the right-hand side (prograde) from reaching the left-hand side of the diagram. The same situation is repeated more or less in the other cases. For this reason, we often read that some collisional event may be at the origin of Venus’ retrograde rotation, the most realistic being related to the transfer of negative angular momentum from a retrograde satellite (see e.g. McCord, 1968; Makarov and Goldin, 2023).
However, the impossibility mentioned above results from considering a frozen model. The inspection of Fig. 9 shows that, if we accept that the physical conditions are certainly not frozen, and the relative contribution of the tidal and atmospheric torques may have been different in the past, it is possible to have paths leading smoothly from a fast prograde rotation in the past to an asynchronous retrograde rotation in the present. We may, for example, consider an evolutive path starting in domain 0 (green area) of Fig. 9 and crossing the frontier between it and domain I. In domain 0, the rotation will evolve towards the synchronous solution and will reach it if enough time is allowed for the path segment inside domain O. When crossing the boundary towards domain I, the synchronous solution becomes unstable, and the rotation is forced to evolve towards one of the two new stable asynchronous solutions. If the rotation is synchronous before the crossing, there is no mathematical reason to privilege one of the two and we have equal probabilities of evolving towards $\mathsf{P}$ or $\mathsf{P}^{\prime}$. However, the increase in the moment of inertia due to the transfer of mass from the planet core to the atmosphere may create a bias in favor of retrograde rotations.
If the atmosphere of Venus is still changing, the current rotational state of Venus can change. For instance, if we assume that Venus’ rotation is currently trapped in a retrograde asynchronous stationary solution, we get $b=0.97489$, which is very close to the boundary $b=1$. A small change, say, in the surface temperature of Venus, may be such that $b$ increases to $b>1$. Hence, stable asynchronous stationary solutions cease to exist and the rotation of Venus will grow continuously.
In fact, we cannot a priori exclude the possibility that the rotation of Venus is already in such a state of runaway growth. The measurements of Venus’ rotation period are difficult, and the available ones (see Margot et al. 2021) do not allow for the exclusion of a very slow growth.
4.1 Smooth transitions in a toy model
We rewrite Eqn. (1) with the modified dependent variable
$y=\frac{\nu}{\gamma^{\prime}}$ (12)
and the positive coefficient
$D=\frac{4A\Gamma}{\gamma^{\prime}}.$ (13)
Hence,
$\dot{y}=D\ \frac{(b-1)y^{3}+(b\Gamma^{2}-1)y}{(y^{2}+1)(y^{2}+\Gamma^{2})}$ (14)
Figure 10: : Left Solution of Eqn. (14) along a path defined by $b=\Gamma/64$ crossing the regions 0 and I of Fig. 9 before reaching the boundary $b=1$. Right The same with a logarithmic scale. Also shown: The locus of the stationary root $y_{0}=\nu_{0}/\gamma^{\prime}$ (dashed line), and the derivative $dy/dt$ (black).
Fig. 10 shows a solution for this equation along a path defined by $b=\Gamma/64$ with initial conditions $y(0)=3$ and $D=10$. This path is arbitrary, starts at the origin $b=0,\Gamma=0$ (absence of atmosphere), and evolves linearly as the atmosphere is formed. It crosses the boundary of domains 0 and I when $\Gamma=4$ and reaches the domain $b\geq 1$ when $\Gamma=64$. The following features may be seen: While $\Gamma<4$, $y$ decreases tending to the synchronous solution $y=0$, which is not reached; When $\Gamma=4$, the solution $y=0$ becomes unstable, and $y$ evolves away from it. In the first segment, the growth is exponential (we have $\dot{y}\simeq y$). It approaches the stationary solution $\mathsf{P}$. Since the parameters $b$ and $\Gamma$ are continuously changing, the position of this solution changes (see the dashed line). It is not reached exactly. The actual solution oscillates around it. Nothing is seen when the boundary $b=1$ is crossed. The rotation velocity continues to grow at the same pace as before.
If the initial condition is taken as negative, the result is the same upside down. Equation (14) can be written as $\dot{y}=yF(y,t)$ , and it is easy to see that no solution can cross the line $y=0$. If at a given moment $y=0$, it will remain equal to 0 forever.
The toy model can be improved by taking into account that the atmosphere is formed by the planet’s outgassing. So, mass is being transferred from the inner parts of the planet to the atmosphere. This process increases the moment of inertia of the system. The effect of a small uniform increase in the moment of inertia of the planet is one small subtracting quantity Z-Z in $dy/dt$. Hence,
$\dot{y}=D\ \frac{(b-1)y^{3}+(b\Gamma^{2}-1)y}{(y^{2}+1)(y^{2}+\Gamma^{2})}-Z.$ (15)
where $Z$ is proportional to the rate of deceleration of the planet.
Fig. 11 shows the initial part of one solution for this equation obtained with the same parameters and initial conditions as before for $Z=10^{-3}$. In this solution, $y$ becomes negative as soon as the solution approaches the synchronous stationary solution, still inside the domain 0. The rotation becomes sub-synchronous (i.e. $\nu<0$ , or $\Omega<n$) and later on becomes retrograde. We stress the fact that what is being observed in Fig. 11 is not the transition from prograde to retrograde, but the transition from super-synchronous ($\Omega>n$) to sub-synchronous ($\Omega>n$). But as $|\nu|$ increases, it can become larger than $2n$ and, hence, $\Omega<0$.
Figure 11: Solution of Eqn. (15) along the same path as the solution in the previous figure, when the moment of inertia growth is responsible for a deceleration $Z=-10^{-3}$. At $t=0\ (\Gamma=0)$, the rotation is prograde. Near synchronization, it changes from supersynchronous ($y>0$) to subsynchronous ($y<0$). Subsequently, it becomes retrograde.
5 Conclusion
The main conclusion of this report is that no collision with other bodies is necessary to convert the rotation of an Earth or super-Earth with an important atmosphere formed in the course of its evolution to retrograde. It is enough for the planet to be at a distance from the host star short enough to allow the tidal torques to almost synchronize the rotation of the planet before the formation of the bulk of its atmosphere (but not so close that the formed atmosphere is destroyed by the stellar radiation). As the atmosphere is formed from the outgassing of the planet, the transfer of mass from the planet to its atmosphere may be responsible for a loss of angular velocity, so that the rotation is subsynchronous when the atmospheric torques become more important than the tidal torques  [^5] Keep in mind that subsyncronous means $\nu<0$ and retrograde means $\Omega<0$. In the continuation, the prevailing atmospheric torques increase the relative rotation of the planet and create a runaway increase of its value, allowing the rotation to become retrograde later on. None of these processes involves a catastrophic event (even if the transformation of the stable stationary solution into an unstable one is one typical example of René Thom’s catastrophes – see Arnold, 1984). The formation of the planetary atmosphere is a continuous, smooth process, which may be more or less efficient but is not a low-probability event.
The example constructed using a toy model in Sec. 4.1 introduced one evolution history of the planet in which the parameters $b$ and $\Gamma$ grow uniformly. This choice is only justified by its simplicity. The mathematical phenomena shown in that example do not depend on the adopted path. The essential points are that the system must remain in domain 0 long time so that it reaches an almost synchronization and then crosses the boundary from domain 0 to domain I where the nature of the synchronous solution changes from stable to unstable. The change to retrograde is not a necessary fate. It depends on the sign of $\nu$ just before crossing. If the motion is still super-synchronous, it will remain prograde forever. Conversely, if it is subsynchronous, it will evolve towards a retrograde rotation.
The dense atmosphere of the planet plays a role not only in the inversion of the planetary rotation, but also in keeping it retrograde. The study of a naked Venus shows that the current retrograde rotation would become prograde again in the absence of the atmospheric torque in less than one million years. The full 3D creep tide theory indicates that, at this inversion, we do not have a simple reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the inversion is completed. The same simulations show that the equinox precession of Venus is not a circulation like the Earth’s, but rather an oscillation.
[…]
Acknowledgements This investigation was sponsored by CNPq (Proc. 303540/2020-6) and FAPESP (Proc. 2016/13750-6 ref. PLATO mission). Part of this research was conducted at the Southwest Research Institute in Boulder.